# Constructing embedded surfaces for cellular embeddings of leveled spatial graphs

### v1

@@ 4.1 Placing cylinders under fragments
Let $\mathcal{G}$ be a leveled embedding in representative leveling with $n$ levels, consecutively embedded in a surface $\mathcal{S}_{i}$. Let $P_{i}$ be a cylinder which is glued to two not necessarily distinct disks $D_{1},D_{2}$, such that $\mathcal{S}_{i}=\mathcal{S}_{i-1}\cup_{\varphi_{i}}P_{i}$. A subset of the spine list of $\mathcal{G}$ consisting of the endpoints of the fragments embedded in $P_{i}$ which lie on one end of $P_{i}$ is called a spine sublist of $P_{i}$. Reversing the order of the endpoints in one of the spine sublist results in the disk sublists of $P_{i}$. This is, to obtain the disk sublists, the boundary components of the cylinder are considered as naturally identified.
@@
If there exists a pair of fragments with endpoints on both ends of $P_{i}$ with two or more halftwists with respect to $P_{i}$, then the two fragments need to cross over and under each other. Therefore, the embedding would not be leveled. Moreover, if $f$ has three or more halftwists with respect to $P_{i}$, then any other fragment with endpoints on both ends of $P_{i}$ needs to be parallel to it on $P_{i}$. Therefore, the difference of halftwists with respect to $P_{i}$ of the two fragments is at most 1. This is impossible due to the previous argument. ∎
@@
    $(s_{1},\dots,s_{n},\alpha_{1},\dots\alpha_{k},\beta_{1},\dots,\beta_{h},s_{n},\dots,s_{1}/t_{1},\dots,t_{q},\beta_{1},\dots,\beta_{h},\alpha_{1},\dots,\alpha_{k},t_{q},\dots,t_{1}),$
@@
    $(a_{1},\dots,a_{m},\alpha_{1},\dots,\alpha_{k},b_{1},\dots,b_{p},\beta_{1},\dots,\beta_{h},a_{m},\dots,a_{1}/\beta_{1},\dots,\beta_{h},\alpha_{1},\dots,\alpha_{k},b_{1},\dots,b_{p});$
@@
For the first part, consider a cylinder $P$ that is not attached to a surface. There are two ways to embed edges in $P$ such that their endpoints lie on the boundary of the cylinder: either the two endpoints of an edge are attached to both ends of the cylinder, or both endpoints are attached to the same end. The endpoints of the fragments embedded in $P_{i}$ can be moved along the connected component of the boundary of $P_{i}$ to which the endpoints belong to while maintaining their embedding in $P_{i}$. Lemma 2 gives conditions for the embedding to be leveled: in order to obtain the lists presented, consider the cylinder $P_{i}$ embedded in $\mathbb{R}^{3}$ but not glued to $\mathcal{S}_{i}$ depicted in fig. 2. Shifting the endpoints of the fragments along the boundary components of $P_{i}$ while respecting the conditions given by lemma 2 yields the six presented lists. ∎
@@
Proposition 2 allows to extend proposition 1:
@@
The neighbourhood of the spine of $\mathcal{G}$ in the surface $\mathcal{S}$ has two components. We call them $N_{u}$ and $N_{l}$. For an endpoint $p$ of a fragment $f$ embedded in a cylinder $P$, let $N_{p}$ be a small neighbourhood of $p$ in $P$. Then $N_{p}$ is contained in exactly one of $N_{u}$ or $N_{l}$. We say that $p$ is an upper endpoint (resp. lower endpoint) if $N_{p}$ is contained in $N_{u}$ (resp. $N_{l}$). (fig. 3)
@@
Consider a cylinder $P_{i}$ of the consecutively embedding of $\mathcal{G}$ in $\mathcal{S}.$
@@
We only prove the case where the disk sublists restricted to $a,b$ and the endpoint of $x$ lying on the spine between the pair of endpoints of $(a,b)$ as in the statement is of the form $(a,x,b/a,b)$. The other cases are similar. We prove the counterpositive: if $x$ is embedded in an upper regular cylinder, then $x$ is at higher level than $a$. (fig. 5)
@@
We only need to prove 2, because 1 follows from proposition 3 and 3 follows from 2 by applying a reflection with respect to the plane where the spine lies. We prove the counterpositive: if $P$ and $P^{\prime}$ are upper cylinders and the endpoints of $a,b,x$ are lower endpoints, then $x$ is at lower level than $a$ (fig. 6). Let $P$ be an upper cylinder and $x$ have a lower endpoint in between two lower endpoints of $a$ and $b$ in the same disk sublist. If $x$ can also be embedded in $P$, then $x$ is at lower level than $b$ due to lemma 3. If $x$ is embedded in another upper consonant cylinder $P^{\prime}$, then it conflicts with either both or exactly one of $a$ and $b$. In the latter case, the crossing number of $x$ with the fragment it does not conflict with must be 0. If $x$ conflicts with $a$, then $x$ is at lower level than $a$, because else $x$ would need to pierce the cylinder $P$ or cross over $P$, making it cross also fragment $b$; contradicting $x$ having crossing number 0 with $b$.
If $x$ conflicts only with $b$, it is not possible that $x$ is at higher level than $b$, because then $P^{\prime}$ would cross over $P$, which implies that $x$ crosses over $a$ non-trivially. This is a contradiction of the fact that $x$ has crossing number 0 with $a$.
The case where $x$ conflicts with both $a$ and $b$ implies that $x$ is at lower level than $a$. Indeed, it is not possible that $x$ is at a level in between the levels of $a$ and $b$, because $P^{\prime}$ would need to pierce through $P$. If $x$ is at higher level than $b$, then $P^{\prime}$ must cross over $P$. But this is not possible because $x$ has a lower endpoint in between two lower endpoints of $a$ and $b$. Therefore, $x$ needs first to cross under $a$ or $b$ before crossing over $b$, which is a contradiction to being leveled. Hence in this case $x$ is at lower level than $a$. ∎
@@ 4.2.1 Conditions to avoid self-intersections of the surface
In case 1, the upper cylinder $P$ has an upper endpoint $e$ while the lower cylinder $P^{\prime}$ has a lower endpoint $e^{\prime}$. The ends of the two cylinders where $e$ and $e^{\prime}$ lie do not intersect each other as the plane where the spine lies separates them. (fig. 7)
@@
Let $P$ and $P^{\prime}$ be upper cylinders. If the fragments $f_{1},\dots,f_{s}$ of $P$ which conflict with a fragment of $g_{1},\dots,g_{t}$ of $P^{\prime}$ are at lower level than the fragments of $P^{\prime}$ they conflict with, then the lower endpoints of $f_{1},\dots,f_{s}$ do not interact with the upper endpoints of $g_{1},\dots,g_{t}$ of $P^{\prime}$ they conflict with. (See fig. 8 for the counterpositive.)
@@ 4.2.2 Conditions to ensure cellularity
In this section, we investigate which hamiltonian leveled spatial graphs that are consecutively embedded are cellular.
@@
    If $P_{i}$ is glued to two different disks $D_{1},D_{2}$ of the cellular embedding of $\mathcal{S}_{i-1}\setminus\mathcal{G}_{i-1}$, then the resulting embedding of $\mathcal{G}_{i}$ in $\mathcal{S}_{i}$ is cellular.
@@
    If $P_{i}$ is glued to a single disk $D$, then the resulting embedding is cellular if and only if there exists at least one conflicting pair of fragments in $f_{1},\dots f_{t}$ in the conflict graph with respect to the boundary of $D.$
@@
    Consider $P_{i}$ glued to $D$ as a torus $T$ with a marked point $t_{0}$, which coincides with the boundary of $D$. Two loops embedded in a torus form a cellular embedding if and only if their crossing number is greater than 0. In our case, loops are fragments and two fragments can cross each other only in $t_{0}$. Therefore, $((T,t_{o}),P_{i})$ is cellular if and only if there exist at least two fragments $f_{1},f_{2}$ such that their endpoints alternate on the boundary of $D.$
@@
Consider the leveled spatial graph $\mathcal{G}$ given by $(3,1,4,2,1,3,4,2).$ We would like to embed $\mathcal{G}$ consecutively such that all endpoints are upper endpoints, fragment $2$ is embedded in the upper hemisphere of a sphere $\mathcal{S}_{0}$, fragment $1$ is embedded in a lower cylinder $P_{1}$ glued to the two disks formed by fragment $2$ (notice that this merges the two disks of the upper hemisphere into a single disk $D$) and fragments $3$ and $4$ are embedded together in an upper cylinder $P_{2}$ glued to $D$. See fig. 9 for a picture of the embedding.
@@
Then, $\mathcal{S}_{i}\setminus\mathcal{G}_{i}$ has $n-2+t$ disks.
@@
The cellular embedding of $\mathcal{G}_{P_{i}}$ in $P_{i}$ has $t$ disks. Indeed, the first fragment of $\mathcal{G}_{P_{i}}$ which is attached to both ends of $P_{i}$ cuts an annulus into a disk, and every other fragment of $\mathcal{G}_{P_{i}}$ cuts an existing disk in two disks, adding one disk to the total per fragment.
If $P_{i}$ is glued to two different disks of $\mathcal{S}_{i-1}$, then $\mathcal{S}_{i}\setminus\mathcal{G}_{i}$ has $n-2+t$ disks.
If $P_{i}$ is glued to a single disk $D$ and the resulting embedding of $\mathcal{G}_{i}$ in $\mathcal{S}_{i}$ is cellular as seen in lemma 4, then $\mathcal{S}_{i}\setminus\mathcal{G}_{i}$ has $(n-1)+(t-1)$ disks. Indeed, for each end of $P_{i}$, the disk of $P_{i}\setminus\mathcal{G}_{P_{i}}$ that contains $\partial P_{i}$ which is not part of the spine is merged with $D$. Notice that, due to cellularity, the disks merged on the two ends of $P_{i}$ are different. ∎
Given that the embedding is cellular, we adapt the “Face tracing algorithm” of a cellular graph embedding explained in [5] to be able to trace the boundaries of its disks.
@@ 4.3 The algorithm for the surface construction
Let $\mathcal{F}_{-\{c,c+1\}}$ be the set of $N$ fragments of $\mathcal{G}$ that are not in level $c$ and $c+1$. Generate all possible partitions $P_{k},1\leq k\leq\text{e}^{-1}\sum_{m=0}^{\infty}\frac{m^{N}}{m!}$ (Dobińsky’s formula [3]), of $\mathcal{F}_{-\{c,c+1\}}$ into disjoint nonempty subsets. That is, the elements of $P_{k}$ are subsets $P_{k}^{j}$, $1\leq j\leq J(k)$ of fragments, where $J(k)$ is the number of the subsets in $P_{k}$. The fragments in $P_{k}^{j}$ will be candidates for fragments embedded in one cylinder which by abuse of notation we also denote by $P_{k}^{j}$.
@@
    • $P_{i}^{h}$ is an upper (resp. lower) cylinder, and $P_{i}^{j}$ is a lower (resp. upper) one, and there exists a fragment of $P_{i}^{h}$ which has an upper (resp. lower) endpoint in between two upper (resp. lower) endpoints of fragments $g_{1},g_{2}$ of $P_{i}^{j}$ belonging to the same spine sublist of $P_{i}^{j}$ such that there is no path from an endpoint of $g_{1}$ to an endpoint of $g_{2}$ along the boundary of the disk with endpoints of $g_{1},g_{2}$ on its boundary which only crosses over (resp. under) $P_{i}^{h}$ or below (resp. above) the fragments embedded in $\mathcal{S}_{0}$ [^1] To detect if a path from an endpoint of $g_{1}$ to an endpoint of $g_{2}$ along the boundary of the disk $D$ with endpoints of $g_{1},g_{2}$ on its boundary only crosses over a cylinder $P_{i}^{h}$, it is enough to trace the boundary of $D$ with 1 and check if there exists a path from the endpoint of $g_{1}$ to the endpoint of $g_{2}$ which includes at least one fragment embedded on a cylinder $P_{i}^{k}$ with $k>h$ and no fragments embedded on $\mathcal{S}_{l}$ for $l<i$ which conflict with $f$. (proposition 5)
@@
If $\mathcal{P}_{i}$ is discarded, consider $\mathcal{P}_{i+1}$. If $\mathcal{P}_{i+1}$ does not exist, set $c=c+1$. If $c=n-1$, the algorithm ends unsuccessfully.
@@
There are $42^{J(2)}=2^{2}=$ 4 possibilities to label the cylinders of $P_{2}$ as upper or lower cylinders: $\{\{A\}^{U},\{B,C\}^{U}\}$, $\{\{A\}^{U},\{B,C\}^{L}\}$, $\{\{A\}^{L},\{B,C\}^{U}\}$, $\{\{A\}^{L},\{B,C\}^{L}\}$.
@@
$\mathcal{P}_{1}$ =$=\{(A_{3,u}^{1}/A_{3,u}^{1})^{U}\},$ $\mathcal{P}_{2}$ =$=\{(A_{3,u}^{1}/A_{3,l}^{1})^{U}\},$ $\mathcal{P}_{3}$ =$=\{(A_{3,l}^{1}/A_{3,u}^{1})^{U}\},$ $\mathcal{P}_{4}$ =$=\{(A_{3,l}^{1}/A_{3,l}^{1})^{U}\},$
$\mathcal{P}_{5}$ =$=\{(A_{3,u}^{1}/A_{3,u}^{1})^{L}\},$ $\mathcal{P}_{6}$ =$=\{(A_{3,u}^{1}/A_{3,l}^{1})^{L}\},$ $\mathcal{P}_{7}$ =$=\{(A_{3,l}^{1}/A_{3,u}^{1})^{L}\},$ $\mathcal{P}_{8}$ =$=\{(A_{3,l}^{1}/A_{3,l}^{1})^{L}\}.$
@@
If we assume that the spine of the leveled spatial graph is a non-hamiltonian cycle, fragments do not need to be edges, but they all are planar connected graphs by definition. In order to place handles under such spatial fragments, we need to attach $n$-punctured spheres instead of $2$-punctured spheres (i.e. cylinders) to build the surface. Note that an $n$-punctured sphere can be obtained by gluing $n-1$ cylinders together.
@@
Let $\mathcal{G}$ be a leveled spatial graph with spine $C$. Let $e$ be any edge of $C$ and choose a spanning tree $T$ of $\mathcal{G}$ which contains $C-e$. The choice of $T-(C-e)$ is independent of the choice of $e$.
Embed $T\cup e$ in a sphere $\mathcal{S}_{0}$, such that $C$ is the equator of $\mathcal{S}_{0}$. This embedding is cellular, because $T\cup e$ is a connected trivial spatial graph. Contracting all edges in $T-C$ gives a hamiltonian leveled spatial graph $\mathcal{G}^{\prime}$ to which algorithm 1 can be applied. If the algorithm terminates successfully, a cellular embedding $(\mathcal{G}^{\prime},\mathcal{S}^{\prime})$ is obtained. The edge contraction can be reversed by vertex splitting and subdividing edges in $\mathcal{S}^{\prime}$ to obtain a cellular embedding $(\mathcal{G},\mathcal{S})$.
@@
Note that the underlying abstract graph of a leveled embedding consists of at most one non-planar 2-connected component, by definition. For example, the graph $K_{5}\_K_{5}$, obtained by joining two copies of $K_{5}$ by an edge (see fig. 11), does not admit a leveled embedding, because it contains two non-planar 2-connected components. However, if multiple leveled embeddings are attached to each other via trivial spatial graphs, the algorithms presented before can still be applied. We call this generalization of leveled embeddings multi-leveled embeddings:

### v2

@@ 4.1 Placing cylinders under fragments
Let $\mathcal{G}$ be a leveled embedding in representative leveling with $n$ levels, consecutively embedded in a surface $\mathcal{S}_{i}$. Let $P_{i}$ be a cylinder which is glued to two not necessarily distinct disks $D_{1},D_{2}$, such that $\mathcal{S}_{i}=\mathcal{S}_{i-1}\cup_{\varphi_{i}}P_{i}$. A subset of the spine list of $\mathcal{G}$ consisting of the endpoints of the fragments embedded in $P_{i}$ which lie on one end of $P_{i}$ is called a spine sublist of $P_{i}$. Reversing the order of the endpoints in one of the spine sublist results in the disk sublists of $P_{i}$. That is, to obtain the disk sublists, the boundary components of the cylinder are considered as naturally identified.
Definition 4.
Let $f$ be a fragment embedded in a cylinder $P$ such that the endpoints of $f$ lie on both ends of $P$. Consider the rectangle representation of $P$ given by $0,1]\times[0,2\pi]/\sim$, where the point $(t,0)$ is identified with the point $(t,2\pi)$ for $t\in[0,1].$ We say that $f$ has a right (resp. left) halftwist with respect to $P$ if $f$ is ambient isotopic with fixed endpoints to the line joining points $(0,0)$ and $(1,\pi)$ (resp. the line joining points $(0,2\pi)$ and $(1,\pi)$). (Figure 2)
Figure 2: A fragment $f$ with a right halftwist with respect to the cylinder $P$.
@@
If there exists a pair of fragments with endpoints on both ends of $P_{i}$ with two or more halftwists with respect to $P_{i}$, then the two fragments need to cross over and under each other. Therefore, the embedding would not be leveled. Moreover, if $f$ has three or more halftwists with respect to $P_{i}$, then any other fragment with endpoints on both ends of $P_{i}$ needs to be parallel to it on $P_{i}$. Therefore, the difference of halftwists with respect to $P_{i}$ of the two fragments is at most 1. That is impossible due to the previous argument. ∎
@@
    $(s_{1},\dots,s_{n},\alpha_{1},\dots\alpha_{k},\beta_{1},\dots,\beta_{h},s_{n},\dots,s_{1}/t_{1},\dots,t_{q},\beta_{1},\dots,\beta_{h},$
    $\alpha_{1},\dots,\alpha_{k},t_{q},\dots,t_{1}),$
@@
    $(a_{1},\dots,a_{m},\alpha_{1},\dots,\alpha_{k},b_{1},\dots,b_{p},\beta_{1},\dots,\beta_{h},a_{m},\dots,a_{1}/\beta_{1},\dots,\beta_{h},$
    $\alpha_{1},\dots,\alpha_{k},b_{1},\dots,b_{p});$
@@
For the first part, consider a cylinder $P$ that is not attached to a surface. There are two ways to embed edges in $P$ such that their endpoints lie on the boundary of the cylinder: either the two endpoints of an edge are attached to both ends of the cylinder, or both endpoints are attached to the same end. The endpoints of the fragments embedded in $P_{i}$ can be moved along the connected component of the boundary of $P_{i}$ to which the endpoints belong to while maintaining their embedding in $P_{i}$. Lemma 2 gives conditions for the embedding to be leveled: in order to obtain the lists presented, consider the cylinder $P_{i}$ embedded in $\mathbb{R}^{3}$ but not glued to $\mathcal{S}_{i}$ depicted in Figure 3. Shifting the endpoints of the fragments along the boundary components of $P_{i}$ while respecting the conditions given by Lemma 2 yields the six presented lists. ∎
@@
Example 2.
Consider the cylinder $P$ on the right of Figure 3. If we name $a_{1}$ the red fragment with endpoints attached to one end of $P$, $\alpha$ the red fragment with endpoints attached to both ends of $P$, $b_{1},b_{2},b_{3}$ the green fragments with endpoints on both ends of $P$ and $x_{1},x_{2},x_{3}$ the green fragments with endpoints on one end of $P$, then the disk sublists of $P$ would be
$(a_{1},\alpha,b_{1},b_{2},x_{1},x_{1},b_{3},a_{1}/\alpha,x_{2},x_{2},b_{1},b_{2},b_{3},x_{3},x_{3}),$
which correspond to the sublists of type $4$ of Proposition 2.
Proposition 2 allows us to extend Proposition 1:
@@
The neighbourhood of the spine of $\mathcal{G}$ in the surface $\mathcal{S}$ has two components. We call them $N_{u}$ and $N_{l}$. For an endpoint $p$ of a fragment $f$ embedded in a cylinder $P$, let $N_{p}$ be a small neighbourhood of $p$ in $P$. Then $N_{p}$ is contained in exactly one of $N_{u}$ or $N_{l}$. We say that $p$ is an upper endpoint (resp. lower endpoint) if $N_{p}$ is contained in $N_{u}$ (resp. $N_{l}$). (Figure 4)
@@
Consider a cylinder $P_{i}$ of the consecutively embedded spatial graph $\mathcal{G}$ in $\mathcal{S}.$
@@
We only prove the case where the disk sublists restricted to $a,b$ and the endpoint of $x$ lying on the spine between the pair of endpoints of $(a,b)$ as in the statement is of the form $(a,x,b/a,b)$. The other cases are similar. We prove the counterpositive: if $x$ is embedded in an upper cylinder, then $x$ is at higher level than $a$. (Figure 6)
@@
We only need to prove 2, because 1 follows from Proposition 3 and 3 follows from 2 by applying a reflection with respect to the plane where the spine lies. We prove the counterpositive: if $P$ and $P^{\prime}$ are upper cylinders and the endpoints of $a,b,x$ are lower endpoints, then $x$ is at lower level than $a$ (Figure 7). Let $P$ be an upper cylinder and $x$ have a lower endpoint in between two lower endpoints of $a$ and $b$ in the same disk sublist. If $x$ can also be embedded in $P$, then $x$ is at lower level than $b$ due to Lemma 3. If $x$ is embedded in another upper cylinder $P^{\prime}$, then it conflicts with either both or exactly one of $a$ and $b$. In the latter case, the crossing number of $x$ with the fragment it does not conflict with must be 0, where the crossing number is the minimum number of crossing points between two fragments up to ambient isotopy. If $x$ conflicts with $a$, then $x$ is at lower level than $a$, because else $x$ would need to pierce the cylinder $P$ or cross over $P$, making it cross also fragment $b$; contradicting $x$ having crossing number 0 with $b$.
If $x$ conflicts only with $b$, it is not possible that $x$ is at higher level than $b$, because then $P^{\prime}$ would cross over $P$, which implies that $x$ crosses over $a$ non-trivially. That is a contradiction of the fact that $x$ has crossing number 0 with $a$.
The case where $x$ conflicts with both $a$ and $b$ implies that $x$ is at lower level than $a$. Indeed, it is not possible that $x$ is at a level in between the levels of $a$ and $b$, because $P^{\prime}$ would need to pierce through $P$. If $x$ is at higher level than $b$, then $P^{\prime}$ must cross over $P$. But that is not possible because $x$ has a lower endpoint in between two lower endpoints of $a$ and $b$. Therefore, $x$ needs first to cross under $a$ or $b$ before crossing over $b$, which is a contradiction to being leveled. Hence in this case $x$ is at lower level than $a$. ∎
@@ 4.2.1 Conditions to avoid self-intersections of the surface
In case 1, the upper cylinder $P$ has an upper endpoint $e$ while the lower cylinder $P^{\prime}$ has a lower endpoint $e^{\prime}$. The ends of the two cylinders where $e$ and $e^{\prime}$ lie do not intersect each other as the plane where the spine lies separates them. (Figure 8)
@@
Let $P$ and $P^{\prime}$ be upper cylinders. If the fragments $f_{1},\dots,f_{s}$ of $P$ which conflict with a fragment of $g_{1},\dots,g_{t}$ of $P^{\prime}$ are at lower level than the fragments of $P^{\prime}$ they conflict with, then the lower endpoints of $f_{1},\dots,f_{s}$ do not interact with the upper endpoints of $g_{1},\dots,g_{t}$ of $P^{\prime}$ they conflict with. (See Figure 9 for the counterpositive.)
@@ 4.2.2 Conditions to ensure cellularity
In this section, we investigate which hamiltonian leveled spatial graphs that are consecutively embedded are cellular. Recall that the subgraph of a consecutively embedded leveled spatial graph $\mathcal{G}$ embedded in a surface $S_{i}$, built successively by gluing the ends of cylinders to a sphere, is called $\mathcal{G}_{i}$.
@@
    If $P_{i}$ is glued to two different disks $D_{1},D_{2}$ of the cellular embedding of $\mathcal{S}_{i-1}\penalty 10000\ \setminus\penalty 10000\ \mathcal{G}_{i-1}$, then the resulting embedding of $\mathcal{G}_{i}$ in $\mathcal{S}_{i}$ is cellular.
@@
    If $P_{i}$ is glued to a single disk $D$, then the resulting embedding is cellular if and only if there exists at least one conflicting pair of fragments in $f_{1},\dots,f_{t}$ in the conflict graph with respect to the boundary of $D.$
@@
    Consider $P_{i}$ glued to $D$ as a torus $T$ with a marked point $t_{0}$, which coincides with the boundary of $D$. Two loops embedded in a torus form a cellular embedding if and only if their crossing number is greater than 0. In our case, loops are fragments and two fragments can cross each other only in $t_{0}$. Therefore, the embedding of $P_{i}$ in the torus $T$ with marked point $t_{o}$ is cellular if and only if there exist at least two fragments $f_{1},f_{2}$ such that their endpoints alternate on the boundary of $D.$
@@
Consider the leveled spatial graph $\mathcal{G}$ given by $(3,1,4,2,1,3,4,2).$ We would like to embed $\mathcal{G}$ consecutively such that all endpoints are upper endpoints, fragment $2$ is embedded in the upper hemisphere of a sphere $\mathcal{S}_{0}$, fragment $1$ is embedded in a lower cylinder $P_{1}$ glued to the two disks formed by fragment $2$ (notice that this merges the two disks of the upper hemisphere into a single disk $D$) and fragments $3$ and $4$ are embedded together in an upper cylinder $P_{2}$ glued to $D$. See Figure 10 for a picture of the embedding.
@@
Then, $\mathcal{S}_{i}\setminus\mathcal{G}_{i}$ has n2+tn-2+t disks.
@@
The cellular embedding of $\mathcal{G}_{P_{i}}$ in $P_{i}$ has $t$ disks. Indeed, the first fragment of $\mathcal{G}_{P_{i}}$ which is attached to both ends of $P_{i}$ transforms an annulus into a disk, and every other fragment of $\mathcal{G}_{P_{i}}$ transforms an existing disk in two disks, adding one disk to the total per fragment.
If $P_{i}$ is glued to two different disks of $\mathcal{S}_{i-1}$, then $\mathcal{S}_{i}\setminus\mathcal{G}_{i}$ has n2+tn-2+t disks.
If $P_{i}$ is glued to a single disk $D$ and the resulting embedding of $\mathcal{G}_{i}$ in $\mathcal{S}_{i}$ is cellular as seen in Lemma 4, then $\mathcal{S}_{i}\setminus\mathcal{G}_{i}$ has (n1)+(t1)(n-1)+(t-1) disks. Indeed, for each end of $P_{i}$, the disk of $P_{i}\setminus\mathcal{G}_{P_{i}}$ that contains $\partial P_{i}$ which is not part of the spine is merged with $D$. Notice that, due to cellularity, the disks merged on the two ends of $P_{i}$ are different. ∎
Given that the embedding is cellular, we adapt the “Face tracing algorithm” of a cellular graph embedding explained in GT (01) to be able to trace the boundaries of its disks.
@@ 4.3 The algorithm for the surface construction
Let $\mathcal{F}_{-\{c,c+1\}}$ be the set of $N$ fragments of $\mathcal{G}$ that are not in level $c$ and $c+1$. Generate all possible partitions $P_{k},1\leq k\leq\text{e}^{-1}\sum_{m=0}^{\infty}\frac{m^{N}}{m!}$ (Dobińsky’s formula Dob (77)), of $\mathcal{F}_{-\{c,c+1\}}$ into disjoint nonempty subsets. That is, the elements of $P_{k}$ are subsets $P_{k}^{j}$, $1\leq j\leq J(k)$ of fragments, where $J(k)$ is the number of the subsets in $P_{k}$. The fragments in $P_{k}^{j}$ will be candidates for fragments embedded in one cylinder which by abuse of notation we also denote by $P_{k}^{j}$.
@@
    • $P_{i}^{h}$ is an upper (resp. lower) cylinder, and $P_{i}^{j}$ is a lower (resp. upper) one, and there exists a fragment of $P_{i}^{h}$ which has an upper (resp. lower) endpoint in between two upper (resp. lower) endpoints of fragments $g_{1},g_{2}$ of $P_{i}^{j}$ belonging to the same spine sublist of $P_{i}^{j}$ such that there is no path from an endpoint of $g_{1}$ to an endpoint of $g_{2}$ along the boundary of the disk with endpoints of $g_{1},g_{2}$ on its boundary which only crosses over (resp. under) $P_{i}^{h}$ or below (resp. above) the fragments embedded in $\mathcal{S}_{0}$.  [^1] To detect if a path from an endpoint of $g_{1}$ to an endpoint of $g_{2}$ along the boundary of the disk $D$ with endpoints of $g_{1},g_{2}$ on its boundary only crosses over a cylinder $P_{i}^{h}$, it is enough to trace the boundary of $D$ with 1 and check if there exists a path from the endpoint of $g_{1}$ to the endpoint of $g_{2}$ which includes at least one fragment embedded on a cylinder $P_{i}^{k}$ with $k>h$ and no fragments embedded on $\mathcal{S}_{l}$ for $l<i$ which conflict with $f$. (Proposition 5)
@@
If $\mathcal{P}_{i}$ is discarded, consider $\mathcal{P}_{i+1}$. If $\mathcal{P}_{i+1}$ does not exist, set $c=c+1$. If c=n1c=n-1, the algorithm ends unsuccessfully.
@@
There are $2^{J(2)}=2^{2}=4$ possibilities to label the cylinders of $P_{2}$ as upper or lower cylinders: $\{\{A\}^{U},\{B,C\}^{U}\}$, $\{\{A\}^{U},\{B,C\}^{L}\}$, $\{\{A\}^{L},\{B,C\}^{U}\}$, $\{\{A\}^{L},\{B,C\}^{L}\}$.
@@
$\mathcal{P}_{1}$ $=\{(A_{3,u}^{1}/A_{3,u}^{1})^{U}\},$ $\mathcal{P}_{2}$ $=\{(A_{3,u}^{1}/A_{3,l}^{1})^{U}\},$ $\mathcal{P}_{3}$ $=\{(A_{3,l}^{1}/A_{3,u}^{1})^{U}\},$
$\mathcal{P}_{4}$ $=\{(A_{3,l}^{1}/A_{3,l}^{1})^{U}\},$ $\mathcal{P}_{5}$ $=\{(A_{3,u}^{1}/A_{3,u}^{1})^{L}\},$ $\mathcal{P}_{6}$ $=\{(A_{3,u}^{1}/A_{3,l}^{1})^{L}\},$
$\mathcal{P}_{7}$ $=\{(A_{3,l}^{1}/A_{3,u}^{1})^{L}\},$ $\mathcal{P}_{8}$ $=\{(A_{3,l}^{1}/A_{3,l}^{1})^{L}\}.$
@@
If we assume that the spine of the leveled spatial graph is a non-hamiltonian cycle, fragments do not need to be edges, but they all are planar connected graphs by definition. In order to place handles under such spatial fragments, we need to attach $n$-punctured spheres instead of $2$-punctured spheres (i.e. cylinders) to build the surface. Note that an $n$-punctured sphere can be obtained by gluing n1n-1 cylinders together.
@@
Let $\mathcal{G}$ be a leveled spatial graph with spine $C$. Let $e$ be any edge of $C$ and choose a spanning tree $T$ of $\mathcal{G}$ which contains CeC-e. The choice of T(Ce)T-(C-e) is independent of the choice of $e$.
Embed $T\cup e$ in a sphere $\mathcal{S}_{0}$, such that $C$ is the equator of $\mathcal{S}_{0}$. This embedding is cellular, because $T\cup e$ is a connected trivial spatial graph. Contracting all edges in TCT-C gives a hamiltonian leveled spatial graph $\mathcal{G}^{\prime}$ to which Algorithm 1 can be applied. If the algorithm terminates successfully, a cellular embedding $(\mathcal{G}^{\prime},\mathcal{S}^{\prime})$ is obtained. The edge contraction can be reversed by vertex splitting and subdividing edges in $\mathcal{S}^{\prime}$ to obtain a cellular embedding $(\mathcal{G},\mathcal{S})$.
@@
Note that the underlying abstract graph of a leveled embedding consists of at most one non-planar 2-connected component, by definition. For example, the graph $K_{5}\_K_{5}$, obtained by joining two copies of $K_{5}$ by an edge (see Figure 13), does not admit a leveled embedding, because it contains two non-planar 2-connected components. However, if multiple leveled embeddings are attached to each other via trivial spatial graphs, the algorithms presented before can still be applied. We call this generalization of leveled embeddings multi-leveled embeddings: